OpenStudy (anonymous):

evaluate log1/2 (1/16)^4 .z.

OpenStudy (anonymous):

@jdoe0001 @agent0smith @Compassionate @kropot72 @thomaster do any of you know

OpenStudy (raden):

solve for x : (1/2)^x = (1/16)^4

OpenStudy (anonymous):

thats not one of the options

OpenStudy (raden):

i mean, assumed the answer is x. :) can you solve x in the equation above ?

OpenStudy (anonymous):

@RadEn tbh.. i just want to know the answer

OpenStudy (raden):

well, hoped this makes sense for you. (1/2)^x = (1/16)^4 (1/2)^x = (1/2^4)^4 (1/2)^x = ((1/2)^4)^4 (1/2)^x = (1/2)^16 identically both exponents, get x = 16